OpenStudy (anonymous):

The equation of the circle whose center is (0, 0) and radius is length 9 is x² + y² = 3 x² + y² = 9 x² + y² = 81

OpenStudy (shamil98):

The formula for a circle is: (x-h)^2 + (y - k)^2 = r^2 (h,k) = center coordinates What do you think it is then?

OpenStudy (shamil98):

and r = radius

OpenStudy (anonymous):

i mean lol I think it should be number 1

OpenStudy (shamil98):

No. Your radius is 9 if the equation is equal r^2 9^2 is?

OpenStudy (anonymous):

I dont know why but im doing a lesson online and it says 9 is wrong ?

OpenStudy (shamil98):

It's not 9 -.-

OpenStudy (shamil98):

(x-0)^2 + (y-0)^2 = 9^2 x^2 + y^2 = 81

OpenStudy (anonymous):

??? I guess not i alredy tried it 2 and it was wrong ?

OpenStudy (shamil98):

\[\large (x-h)^2 + (y-k)^2 = r^2\] \[\large h = 0 , k = 0 , r= 9\] \[\large x^2 + y^2 = 81\]

OpenStudy (anonymous):

I think it could be right thanks.